Mole fraction of the solute in a 1.00 molal aqueous solution is
(1) 1.7700
(2) 0.1770
(3) 0.0177
(4) 0.0344

Grade:upto college level

5 Answers

Deepak Patra
9 years ago
Sol. Mole fraction of solute = 1/56.55 = 0.0177 Answer : (3)
Rajveer Singh Rathore
6 years ago
1.00 molal aqueous solution means that 1 mol of solute is dissolved in 1 Kg of water.Hence,No. Of moles of solute = 1No. Of moles of water = 1000/18=55.556MOLE FRACTION OF SOLUTE = No. Of moles of solute/no. Of moles of solute + No. Of moles of solventMOLE FRACTION = 1/1+55.556 =0.1768[0.0177(approx.) :):):)
Poorvee
5 years ago
Mole of water is 1000/18 =55.55 mole
Mole of solute is 1/1= 1 mole then
Mole friction of solute = 1/1+55.55 =0.1768 =0.177
 
Shahensa kobir
5 years ago
Moles of water=
mass/molecular mass=1000g/18g=55.55moles.
Mole fraction of solute=(no of moles of solute)/(no of mole of solute+no of moles of solvent)=1/(1+55.55)=0.177
Sahithi pendyala
5 years ago
1 molal of aqueous solution mean 1 MOLE OF SOLUTE PER 1KG SOLVENT.                        1 mole of water for 18 gms.   55.5 moles of water per 1000 gms.                                                Mole fraction is ratio of moles of solute to total moles of solution.  Xa÷(Xa+Xb)                                         1 (1+55.5)=0.0177